Minutes of the Tracker Technical Meeting

November 6, 2003



      1. Action items from last week

      2. Monday delta-PRR preparations

      3. Sidewall, coupon tests, static sidewall, fixtures, and other vibe-test prep status

      4. Torque striping

      5. T/V prep status; TRR date; dummy TEM issues

      6. A.O.B.


New Action Items:

  1. Tom: discuss the new vibration test information from Italy with NASA.
  2. Robert: request confirmation from Gunther about EGSE hardware purchases.
  3. Sandro: send document drafts 1601 and 1918, as well as the update to 1584.
  4. Sandro: complete inspection of the heavy cores.
  5. Jeff: send the airbill numbers to Sandro for the vibration fixture shipments.
  6. Erik: send the vibration predictions to Bari in Excel format.
  7. Sandro: send to Tom information on torque striping paint that they have available.
  8. Sandro: propose a detailed schedule for the T/V test.


Old Actions:

  1. Nanda: make up an agenda for the mid-tray follow-up PRR.   (Closed)
  2. Tom: schedule a conference for this follow-up PRR (Closed)
  3. Tom, BJ: finalize and release drawings for the bottom-tray closeouts.  Keep COI stopped until this is ready.  (Still open; in progress)
  4. Tom: estimate delivery times for the titanium parts and bottom-tray closeouts.   (Still open: Jerry is working on moving the Ti part procurement forward to get bids from three vendors.)
  5. Sandro: write down a plan for integrating the EM tower into the T/V chamber and test setup.  (Still open)
  6. Jack et al: write down thermocouple locations, thermal blanket specification, TEM temperature control requirements, dwell time requirements, and T/V test cases.  (Sandro needs much of this by next Tuesday!) (Closed)
  7. Bari: update the T/V test plan and procedure. (Still open)
  8. Tom: write out a revised plan for getting to the vibe test, given fire delays at COI (Closed, but schedule still tight)
  9. Tom: update drawing notes for the sidewalls regarding the flatness requirement (Closed by BJ)
  10. Tom: send COI coupon test data to Italy (Closed)
  11. Sandro: find and send to SLAC a precise completion date for the Plyform coupon tests (Closed; see below)
  12. Luca: check on how many old Motorola VME boards can be had off the shelf in Italy. (Closed; see below)



Sandro reported on the vibration test plans.  See the minutes below from a meeting held yesterday at Alenia.  Some details of the sine test still are at issue.  Sandro estimates that we need 1.5 days per axis, which increases the test time.  Tom will take vibration info just received and talk to NASA about it.  Tom said that he added a note to the procedure regarding Chris Fransen’s email to use limiting during the vibration if necessary to avoid over testing.  The program will pause after the –6dB point to evaluate the results and make sure that the further testing will not induce loads higher than expected.


There was some discussion of the T/V test planning.  See the minutes below from the meeting at Alenia.  Tom said that long cycle and plateau times in the original plan are not necessary for the engineering model.  Jack and Jeff has prepared a few slides (attached) on their test proposal.  The following two paragraphs relate to those slides:


      1.  The exact profile of the test cycles.  The tracker power levels will not vary much, 8.7 - 10.0 W; values are taken from LAT-TD-0225, "Summary of LAT Dissipated Powers".  Concerning the cold plate temperature, given that the tracker runs about 9 - 10C above the GRID temperature (for the HOT case), and predictions (from the CDR) give a tracker temperature of 29C, then for these math model correlation tests, the cold plate must operate to at least 20C; 25C would be better.  At the cold limit, the plate must operate down to about 10C (this number may be smaller than the hot case) below the cold operating limit; i.e., -15C.  For thermal cycling, to achieve qualification levels if that is what you are planning to do, the cold plate must achieve -30 to +50C.


      2.  Definition of thermal balance.  There is of course no exact definition, but for this test article, wherein thermal masses can vary widely, after Jeff makes several transient runs with his thermal model of just the tracker, he will be able to identify those masses which time-lag the most (he probably already has good idea).  The rate of change of those identified temperatures should be no more than 0.1 to 0.2C per hour for about 4 hours, but additionally, the second rate of change (d2T/dt2) should be decreasing during that time period.


Sandro expressed concern that there is not enough time for the test to assure completion before the holiday shutdown at Alenia on Dec 19.  He prefers to prepare for the test in December and to start the T/V cycles on Jan 7.  Sandro asked about requirements for space between the tower and the thermal blanket.  About 1cm was deemed to be okay.


Tom said that the TEM cables will be in contact with cold plate, so heat will shunt into the plate there.  We will only need to monitor the temperature of TEM.  Sandro said that the TEM will have a blanket.  Nanda commented that for real TEM it will be best to keep it at constant temperature throughout the program, to avoid risking delays caused by TEM malfunction.


Luca reported a negative response on obtaining the old VME processor boards in Italy.  Anyway, Gunther managed to get the boards.  Pisa will buy crates and power supplies.  He needs a reply from Gunther in order to send out an order.  Pisa needs to spend the money by the end of the month.  They will also buy the computers.  What else can they buy?  There was a question about how many spare VME processors will be available.


Sandro said that he will send the missing documents (1601 and 1918) and updated document 1584 tomorrow.  Mid tray drawings are being finalized by BJ.  The assembly drawing was sent out yesterday.  (See http://www-glast.slac.stanford.edu/Tracker-Hardware/readiness/Tray_Panel/Tray_Panel_PRR.html for the latest review material.)  There remains some TBD regarding the black paint.  Sandro said that he intends to use the nonconductive variety.  BJ said that he took care of Jerry's comments.  There is still an issue of primer and thinner for the black paint.  Luisella is contacting the manufacturer for application information.  It was agreed that we should not proceed without the correct primer and thinner.  We need to find the delivery time for primer and thinner.  Tom estimates two weeks.  Plyform won't start production without having all materials in hand.  Thick tungsten is still being cut and will arrive at the end of this month.  Also, the bias circuits won’t be available until end of month; they should be available to ship on Nov 24.


Pisa has inspected all the light honeycomb, for a total of 276 pieces.  Sandro has sent the statistics.  Pisa will now inspect the heavy material.  They expect less problems with it.  Sandro needs to know if they can straighten cells manually (the answer so far is NO).  Some cores have some cell walls bent over, probably during shipping and handling.  Plyform will start production with perfect cores.  Sandro is waiting for disposition of the others.


Coupons have their inserts installed today at COI, so they will soon begin pulling.  COI is completing the static sidewalls.  The static-sidewall insert installation tooling arrived.  More M4 inserts were sent yesterday from the newly arrived flight lot.  COI will ship the static sidewalls Saturday morning to Hytec, to arrive Monday for final prep and the static test.  Sandro has not received the vibration fixtures.  Jeff needs to send the airbill numbers.  More than a week has passed since they were shipped. 


Bari would like the predictions data send to them in Excel format, to use for over-plotting the test data.  Erik said that he could send that information.


Plyform is machining sidewalls.  Beginning of next week they will install inserts. Tom noted that only one set of mini sidewalls is needed from Plyform.  The will cut the coupons next week and put in inserts the end of next week.  Sandro will request that they complete the testing without inserts by the end of next week.  Testing of Plyform coupons with inserts will be done at COI.  COI-made coupons will be pulled Monday.  All data should be available by Wednesday evening.  It was clear that the Plyform coupons with inserts will not be pulled, according to this plan, until after the vibration test.  There was some discussion of the risk involved in this.  The decision was to move forward with the test based on the pull tests of non-insert coupons plus the insert coupon and mini-sidewall tests done with COI parts.


Tom reminded us that a torque stripe needs to be on each screw on the sidewall.  Standard "torque striping paint" should be used.  Sandro will send the name of what they have available.


Getting back to the T/V test, Sandro confirmed that it cannot span the Christmas break.  He estimated that the time to turn the tower around after the vibration test is at minimum one week.  He would like more time to test the instruments at Pisa before going to Alenia.  Sandro will propose a schedule for the T/V test.  He doesn't like to leave the tower in Alenia over the break.




Nov, 5th 2003 Alenia meeting minutes


**** EM Vib test.

The Alenia people request a different time schedule in order to test 1

axis in 1.5 day: this is due to their expierence in this kind of activity,

taking also into account the data analysis and discussion and eventually

further actions. They also confirm the 2 hours/strain channel: this time

is requested to glue the strain, to plug, to check and to calibrate it to

the DAQ. So, the effective schedule should be the following:


1st day (Mon Nov, 17th): Incoming tower + instrumentation

2nd day (Tue Nov, 18th): Instrumentation set-up continued

3rd day (Wed Nov, 19th): Z axis

4th day (Thu Nov, 20th): Z axis completion and shaker rotation

5th day (Fri Nov, 21st): X axis

6th day (Mon Nov, 24th): X axis completion and Y axis

7th day (Tue Nov, 25th): Y axis

8th day (Wed Nov, 26th): Test completion and dismounting


For the sinusoidal vibration test at low frequency it can be performed as



Lateral (X&Y) 5   - 7.39 Hz at 20 mm 0-peak

              7.39 - 15  Hz at 4.4 g full level


Thrust (Z)    5   - 5.57 Hz at 20 mm 0-peak

              5.57 - 18  Hz at 2.5 g full level


Of course the 1/4 level test can be performed without any change. 


The additional Alenia request for this tets in order to change the sweep

rate in the 25-35 Hz range is that the test must be divided into three

sub-tests: the first one in the 5-25 Hz at 2 oct/min, the second one in

the 25-35 Hz at 0.75 oct/min, finally the last one up to 50 Hz at 2



For the sine burst test: the Alenia people are evaluating the possibility

to perform the test as you asked. On the contrary, a sort of sine sweep

will be perfomed: in this case they are evaluating the sweep rate in order

to perform a given number of cycles.


Alenia will send us the list of strain gages available in the

laboratory, then we'll send you this list expecting that you give us

a priority list of the strain type, number and location of the strain

gages to be used. Then on the basis of the time allocated for the

instrumentation, Alenia people will try to set-up as much strain gages as

possible to fit the schedule.  



Minutes from the meeting at Alenia, from Nicola Mazziotta:


** EM T/V test


MLI blanket: Alenia can provide it but they need the exact dimensions. The

time needed to have it is 1- 2 weeks.


A tentative schedule proposed by Alenia for the test is:

- 3-4 days for the test set-up

- 9 days of test if the 156 hr total time is confirmed

- 2 days for the completion and dismounting procedure


Some additional days are also needed to release the data in electronic



The Alenia DAQ cannot read the thermistors on the cables, so we can

eventually try to read them with a voltmeter.


We are writing the plan and procedures and the interface plate and the

thermal blanket support are being designed.

The requests are the following:

- which is the exact profile of the cycles?

- thermal balance definition

- minimum distance between the tower and the thermal blanket

- is it possible to interrupt the cycles maintaining the item under vacuum

at room temperature?


Alenia last work day before Christmans is December 19th. On the basis of

their requests the test should begin no later than the end of November.



This is in conflit with other activities:

0) prepare the blanket (drawings, order, procurement)

1) shipping of the tower in Pisa after Viv test (Nov, 27th)

2) dismount the tower

3) check each trays for eventual damages

4) black paint the sidewalls

5) mount the tower again with the new cables and thermocouples

6) ship the tower back to Alenia


The December work period is short due to the Christmas holiday. We can

perform the T/V test starting from January 7th.